A series RLC circuit has a resistor of 25 ohms, a capacitor of 0.05F and and inductor of 1.0 H. What is the impedance of the circuit if it powered by an ac voltage source with an angular frequency of 20 s-1?
Answer #1

Inductive reactance is given by:

XL = w*L

w = angular frequency = 20 s^-1

L = inductance = 1.0 H

Capacitive reactance is given by:

Xc = 1/(w*C)

C = Capacitance = 0.05 F

R = resistance = 25 ohm

Impedance is given by:

Z = sqrt (R^2 + (XL - Xc)^2)

Z = sqrt (25^2 + (20*1.0 - 1/(20*0.05))^2)

Z = 31.4 ohm = Impedance of RLC circuit

Let me know if you've any query.